OpenStudy (anonymous):

What is the exact value of cot60degrees.

OpenStudy (anonymous):

\[\cot=\frac{ 1 }{ \tan }\] \[\tan60=\sqrt{3}\] so, \[\cot60=\frac{ 1 }{ \sqrt{3} }\] \[\cot60=\frac{ 1 }{ 3 }\sqrt{3}\]